OpenStudy (anonymous):

all numbers at least 3 units from -1

OpenStudy (anonymous):

you want this in absolute value notation?

OpenStudy (anonymous):

absolute value inequality

OpenStudy (anonymous):

as inequalities it would be \(x+1\leq 3\) or \(x+1\geq 3\) as absolute value, it is \(|x+1|\geq 3\)